OpenStudy (helpjebal):

How do I simplify this expression? cos^2 x + sin^2 x / cot^2 x - csc^2 x

OpenStudy (helpjebal):

\[\frac{ \cos^2 x + \sin^2 x }{ \cot^2 x - \csc^2 x }\]

OpenStudy (brrandyn):

Recall corresponding identities. \[\cot x = \frac{ \cos }{ \sin }\] & \[\csc x = \frac{ 1 }{ \sin }\] Replace the cotangent and cosecant with thereof. \[\frac{ \cos^2x + \sin^2x }{ \frac{ \cos^2 }{ \sin^2 } - \frac{ 1 }{ \sin^2 } }\] From there, are you able to solve for the solution?

OpenStudy (welshfella):

the numerator = 1

OpenStudy (brrandyn):

^ Yes. It's a Pythagorean identity.

OpenStudy (helpjebal):

so that leaves it as \[\frac{ \cos^2 x + \sin^2 x }{ \frac{ \cos^2 }{ \sin^2 } }\] right? where do I go from there?

OpenStudy (helpjebal):

ohhhh sin^2 x + cos^2 x = 1 is a pythagreon identity but what do we do with the denominator?

OpenStudy (daniel.ohearn1):

1/(cot^2) = tan^2(x)

OpenStudy (daniel.ohearn1):

because cos^2 + sin^2 = 1 and cos^2/sin^2 = cot^2

OpenStudy (helpjebal):

I'm finding this really hard to understand... I'm sorry

OpenStudy (welshfella):

cos^2 x 1 ------ - --- sin^ 2 x sin^2 x = cos^2 x - 1 --------- sin^2 x

OpenStudy (welshfella):

= - sin^2 x -------- = -1 sin^2 x

OpenStudy (welshfella):

so we end up with 1 / -1 = -1

OpenStudy (helpjebal):

how did -sin^2x end up in the numerator?

OpenStudy (welshfella):

cos^2 x + sin^2 x = 1 cos^2 x - 1 = -sin^2 x

OpenStudy (welshfella):

- subtract sin^2 x and 1 from both sides

OpenStudy (helpjebal):

ohhhh I see! thank you!!

OpenStudy (welshfella):

yw